If a random variable X₁ has mean 4 and variance 9, while the random variable X₂
has mean − 2and variance 5, and the two are independent, find
i) E(2X₁ + X₂ - 5)
ii) Var(2X₁ + X₂ - 5)
1
Expert's answer
2019-03-27T06:29:53-0400
"E(X_1)=4, Var(X_1)=9, E(X_2)=-2, Var(X_2)=5."

"i) E(2X_1+X_2-5)=2E(X_1)+E(X_2)-5=2*4-2-5=1"



ii) If X1 and X2 are independent then



"Var(2X_1+X_2-5)=2^2Var(X_1)+Var(X_2)=4*9+5=41"